OpenStudy (anonymous):

HELP PLEASE! The graph of the sine curve below is of electromagnetic energy that represents indigo light: sine graph with points at 0, 0 and 107 and 5 tenths, 1 and 215, 0 and 322 and 5 tenths, negative 1 and 430, 0 What function accurately represents the sine curve for indigo light?

OpenStudy (anonymous):

@barbillus do you think you can help me?

OpenStudy (anonymous):

f(x) = sin 430πx f(x) = sin 215πx f(x) = sin pi over 215x f(x) = sin pi over 430x

OpenStudy (anonymous):

@abb0t @Abmon98 @Ashleyisakitty

OpenStudy (anonymous):

@campbell_st i think its d am i wrong?

OpenStudy (anonymous):

Thanks for helping me! :-)

OpenStudy (campbell_st):

well looking at the data, it takes the curve 430 units to complete 1 cycle... so the period is \[430 = \frac{2\pi}{B}\] then the form of the equation is \[y = Asin(Bx)\] from the data the amplitude is 1 so its \[y = \sin(Bx)\] you need to find the period B using the information shown earlier in the post. hope it helps

OpenStudy (anonymous):

so is it 430? @campbell_st

OpenStudy (campbell_st):

no you need to find B knowing \[430 = \frac{2\pi}{B}\] you need to manipulate this equation to make B the subject

OpenStudy (anonymous):

so its pi/430

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

i mean pi/215?

OpenStudy (campbell_st):

thats correct... so its choice c or the 3rd one.

OpenStudy (anonymous):

yay!! thanks! ;-)